OpenStudy (anonymous):

Solve the volume formula V= lwh for l. A. l=Vwh B. l=V/wh C. l=V/h+w D. l=wh/V

OpenStudy (anonymous):

i=v/wh

OpenStudy (shamim):

Ok, tell me If 8=4*2 Then 4=8/2 Right?

OpenStudy (usukidoll):

Whenever we see "For l" it means that we need l by itself. all we have to do is to divide wh on both sides